A firm has the following demand function 𝑷 = πŸ”πŸŽ βˆ’ 𝟎. πŸ“π‘Έ and its total cost are defined by 𝑻π‘ͺ = πŸπŸ‘ + 𝑸.


a. Find the maximum revenue.


b. Find the production to optimize the profit.


c. Verify if the marginal revenue and marginal cost are the same at the profit-maximizing production level.Β 


1
Expert's answer
2022-04-06T18:02:28-0400

a) Maximum revenue is determined from the equation TR'=0. TR=PQ=(60-0.5Q)Q=60Q-0.5Q2

TR'=60-Q=0; Q=60

b) Optimal profit is determined from the equation MR=MC. MR=TR'=60-Q; MC=TC'=(13+Q)'=1; 60-Q=1; Q=59

c) At Q=59 MR=60-59=1=MC